OpenStudy (anonymous):

WILL MEDAL AND FAN Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 4−x and y = 2x + 3 intersect are the solutions of the equation 4−x = 2x + 3. Part B: Make tables to find the solution to 4−x = 2x + 3. Take the integer values of x between −3 and 3. Part C: How can you solve the equation 4−x = 2x + 3 graphically?

OpenStudy (boldjon):

For B Make a table - I shall hand you the Data.

OpenStudy (boldjon):

4-(1) = 2(1) + 3 4-(2) = 2(2) + 3 4-(3) = 2(3) + 3 4-(0) = 2(0) + 3 4-(-1) = 2(-1) + 3 4-(-2) = 2(-2) + 3 4-(-3) = 2(-3) + 3

OpenStudy (boldjon):

False - 1st one As 3 ≠ 5 Second one as well - As 2 ≠ 7 3rd one as well - As 1 ≠ 9

OpenStudy (boldjon):

4th one as well - As 4 ≠ 3 5th one as well - As 5 ≠ 1 6th one as well - As 6 ≠ -1 Last one as well - As 7 ≠ -3 for c http://www.wolframalpha.com/input/?i=4-x+%3D+2x+%2B+3+